### v1

# Systematical study of exotic $1^{-+}$ tetraquark spectroscopy

Abstract
The masses of exotic quantum-number $1^{-+}$ compact tetraquark states are calculated in a constituent quark model (CQM), where Cornell-like potentials are employed as the central potentials, spin-spin and spin-orbit coupling derived from Breit-Fermi interaction are treated as hyperfine corrections, and model parameters are taken from previous works. The ground state $1^{-+}$ P-wave tetraquarks are predicted at 1.9, 4.2, and 6.6 GeV for the light, charmonium-like, and fully-charm sectors, respectively.
The decay width ratios of $1^{-+}$ tetraquark states are calculated for possible strong two-body decay channels within the rearrangement mechanism, including $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks; $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks; $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks; and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks. The theoretical results are compared with the observed exotic $1^{-+}$ states, and possible channels for observing $1^{-+}$ tetraquark states are also discussed. The work suggests that $\eta_{1}(1855)$ is unlikely to be a compact tetraquark state.
I Introduction
Understanding the spectrum and internal structure of hadrons remains a central goal of quantum chromodynamics (QCD), and states beyond the conventional $q\bar{q}$ mesons and qqqqqq baryons, such as multiquark configurations and hybrids, are now an active frontier of experiment and theory.
Charmonium-like states with total angular momentum $J=1$ have been widely searched over the last two decades and are promising P-wave tetraquark candidates. Several exotic states in the charmonium region, known as Y states with $J^{PC}=1^{--}$ quantum numbers that mix with conventional $\psi$ mesons, could be P-wave tetraquark candidates: Y(4230) Aubert et al. (2005), Y(4360) Wang et al. (2007), Y(4660) Lees et al. (2014), the more recent Y(4484) Ablikim et al. (2022a), and Y(4544) Ablikim et al. (2024).
In conventional mesons composed of a quark and an antiquark $q\bar{q}$, the quantum numbers $J^{PC}$ follow simple rules: parity $P=(-1)^{L+1}$ and charge conjugation $C=(-1)^{L+S}$, where L is orbital angular momentum and S is total spin Navas et al. (2024). These rules allow common states such as $1^{--}$ (the $\rho$ meson) or $2^{++}$ (the $f_{2}$ meson) but not $1^{-+}$. Thus, a $1^{-+}$ state, once observed, is naturally considered to have a nonconventional structure, such as a hybrid meson, hadronic molecule, or tetraquark.
The experimental status of exotic $1^{-+}$ states has advanced through complementary approaches. The existing data provide a foundation for evaluating their properties, which we briefly review below:
The VES (Vertex Spectrometer) experiment observed the exotic $\pi_{1}(1400)$ state ($J^{PC}=1^{-+}$) in $\pi^{-}p\to\pi^{0}\eta n$ decays at 100 GeV Alde et al. (1988), with a mass of $1406\pm 20$ MeV and a width of $180\pm 30$ MeV. The analysis of the $\eta\pi^{0}$ channel showed a clear exotic signal inconsistent with conventional $q\bar{q}$ states.
The $\pi(1600)$ state was first observed by the Brookhaven E852 experiment Adams et al. (1998), which identified an exotic $1^{-+}$ resonance in the $\rho\pi$ channel. The analysis found a mass of $1593\pm 8^{+29}{-47}$ MeV and a width of $168\pm 20^{+150}{-12}$ MeV. Later, E852 Kuhn et al. (2004) observed two exotic $J^{PC}=1^{-+}$ resonances in $\pi^{-}p\to f_{1}(1285)\pi^{-}p$ collisions at 18 GeV: the known $\pi_{1}(1600)$ with mass $M=1709\pm 24\pm 41$ MeV and width $\Gamma=403\pm 80\pm 115$ MeV, and a new $\pi_{1}(2015)$ state at $M=2001\pm 30\pm 92$ MeV with width $\Gamma=333\pm 52\pm 49$ MeV.
In the updated PDG review, $\pi(1400)$ and $\pi(1600)$ are suggested to be a single state, as coupled-channel analyses favor one broad isovector resonance in the 1.4–1.6 GeV region Navas et al. (2024). Of the reported candidates, the $\pi_{1}(2015)$ has only been observed by the E852 experiment Kuhn et al. (2004), and its existence remains inconclusive.
The BESIII Collaboration recently observed the isoscalar exotic state $\eta_{1}(1855)$ with $J^{PC}=1^{-+}$ in $J/\psi\to\gamma\eta\eta^{\prime}$ decays Ablikim et al. (2022b), measuring its mass as $(1855\pm 9^{+6}_{-1})$ MeV, its width as $(188\pm 18^{+3}_{-8})$ MeV with $19\sigma$ significance, and the product branching fraction $\mathcal{B}(J/\psi\to\gamma\eta_{1}(1855))\times\mathcal{B}(\eta_{1}(1855)\to\eta\eta^{\prime})=(2.70\pm 0.41^{+0.16}_{-0.35})\times 10^{-6}$. The narrow width of $\eta_{1}(1855)$ contrasts sharply with the broad $\pi_{1}(1600)$, suggesting different underlying dynamics between these exotic states.
In this work, we study the mass spectrum and decay properties of the $1^{-+}$ P-wave light, charmonium-like, and fully-charm compact tetraquark states within a constituent quark model framework.
The paper is organized as follows. In Sec. II, all possible configurations of the color, spin, flavor, and spatial degrees of freedom of tetraquark states are introduced, and a constituent quark model developed in our previous work Zhao et al. (2021, 2022, 2025) is briefly reviewed. In Sec. III, the $1^{-+}$ P-wave tetraquark mass spectra and decay branching ratios are evaluated within the constituent quark model. The theoretical results are compared with experimental data, and discussions of $\pi_{1}(2015)$ and $\eta_{1}(1855)$ are provided. A summary is given in Sec. IV.
II THEORETICAL MODEL
II.1 Quark configurations of tetraquark
The construction of tetraquark states follows the principle that a tetraquark must be a color singlet, meaning the color wave function must correspond to a $222]_{1}$ singlet within the $SU_{c}(3)$ group. Furthermore, the total wave function of any cluster of identical fermions must be antisymmetric.
In this work, we consider $J^{PC}=1^{-+}$ light ($q_{1}q_{2}\bar{q}_{3}\bar{q}_{4}$), charmonium-like ($q_{1}c_{2}\bar{q}_{3}\bar{c}_{4}$), and fully-charm ($c_{1}c_{2}\bar{c}_{3}\bar{c}_{4}$) tetraquark states, where $q$ and $c$ denote a light quark and a charm quark, and $\bar{q}$ and $\bar{c}$ denote a light antiquark and a charm antiquark, respectively.
The permutation symmetry of the two-quark cluster ($qq$, qcqc, or $cc$) in tetraquark states is described by the Young tableaux $2]_{6}$ and $11]_{\bar{3}}$ of the $SU_{c}(3)$ group, while the color configuration of the two-antiquark cluster ($\bar{q}\bar{q}$, $\bar{q}\bar{c}$, or $\bar{c}\bar{c}$) consists of a $211]_{3}$ triplet and a $22]_{\bar{6}}$ antisextet.
Thus, a $222]_{1}$ color singlet for tetraquark states requires the following color configurations: $2]_{6}\otimes[22]_{\bar{6}}$ and $11]_{\bar{3}}\otimes[211]_{3}$, which correspond to the color sextet-antisextet ($6_{c}\otimes\bar{6}_{c}$) and triplet-antitriplet ($\bar{3}_{c}\otimes 3_{c}$) configurations, respectively.
The $J^{PC}=1^{-+}$ tetraquark states correspond to $L=1$, $S=1$ configurations according to the parity ($P=(-1)^{L}$) and charge-conjugation ($C=(-1)^{L+S}$) quantum numbers, where $L$ is the total orbital angular momentum and $S$ is the total spin of the tetraquark. Thus, the $J^{PC}=1^{-+}$ tetraquark requires an odd $L$, with the lowest possible value being $L=1$. The complete basis is constructed by coupling harmonic-oscillator wave functions, which takes the general form:
$\psi_{NL}$ $=\sum_{{n_{\chi_{i}},l_{\chi_{i}}}}A(n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}})$ (1)
$\times\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\otimes\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\otimes\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3})$
where $\psi_{n_{\chi_{i}}l_{\chi_{i}}}$ are harmonic oscillator wave functions. The sum $n_{\chi_{i}},l_{\chi_{i}}$} runs over $n_{\chi_{1}},n_{\chi_{2}},n_{\chi_{3}},l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$. $N$ and $L$ are, respectively, the total principal quantum number and total angular momentum, where $N=(2n_{\chi_{1}}+l_{\chi_{1}})+(2n_{\chi_{2}}+l_{\chi_{2}})+(2n_{\chi_{3}}+l_{\chi_{3}})$, and $L=l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$ for the P wave tetraquark. The relative Jacobi coordinates $\vec{\chi}_{1}$, $\vec{\chi}_{2}$, and $\vec{\chi}_{3}$ are defined as:
$\vec{\chi}_{1}=\frac{1}{\sqrt{2}}(\vec{r}_{1}-\vec{r}_{2}),$
$\vec{\chi}_{2}=\frac{1}{\sqrt{2}}(\vec{r}_{3}-\vec{r}_{4}),$
$\vec{\chi}_{3}=\frac{m_{1}\vec{r}_{1}+m_{2}\vec{r}_{2}}{m_{1}+m_{2}}-\frac{m_{3}\vec{r}_{3}+m_{4}\vec{r}_{4}}{m_{3}+m_{4}},$ (2)
where $\vec{r}_{j}$ and $m_{j}$ are the coordinate and the mass of the jth quark, respectively. The reduced quark masses are defined as:
$u_{1}=\frac{2m_{1}m_{2}}{m_{1}+m_{2}},$
$u_{2}=\frac{2m_{3}m_{4}}{m_{3}+m_{4}},$
$u_{3}=\frac{(m_{1}+m_{2})+(m_{3}+m_{4})}{m_{1}+m_{2}+m_{3}+m_{4}},$ (3)
The complete bases of the tetraquarks with $N=1$ are listed in Table 1 for reference. The principal quantum number can be increased for each basis to expand the basis size. In this work, basis states with $N\leq 13$ are used to calculate the mass spectrum.
Table 1: The complete bases of tetraquark with quantum number, $N=2n+L$ and $L=1$.
$\Psi_{NL}$ Bases $\psi_{n_{\chi_{1}}l_{\chi_{1}}}(\vec{\chi}_{1}\,)\psi_{n_{\chi_{2}}l_{\chi_{2}}}(\vec{\chi}_{2}\,)\psi_{n_{\chi_{3}}l_{\chi_{3}}}(\vec{\chi}_{3}\,)$
$l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$}
$\Psi_{11}$ {0, 0, 1} $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,1}(\vec{\chi}_{3}\,)$,
{0, 1, 0} $\psi_{0,0}(\vec{\chi}_{1}\,)\psi_{0,1}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$,
{1, 0, 0} $\psi_{0,1}(\vec{\chi}_{1}\,)\psi_{0,0}(\vec{\chi}_{2}\,)\psi_{0,0}(\vec{\chi}_{3}\,)$
For constructing $S=1$ configuration, the possible spin combinations are $\left[\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}\right]_{S=1}$, $\psi_{[s=1]}^{q_{1}q_{2}}\otimes\psi_{[s=0]}^{\bar{q}_{3}\bar{q}_{4}}$, and $\psi_{[s=0]}^{q_{1}q_{2}}\otimes\psi_{[s=1]}^{\bar{q}_{3}\bar{q}_{4}}$.
Because a $qq\bar{q}\bar{q}$ or $cc\bar{c}\bar{c}$ tetraquark must be a color singlet and simultaneously antisymmetric under any permutation of identical quarks, the $qq$-cluster spatial-spin-flavor configurations for $qq\bar{q}\bar{q}$ states are $\psi^{osf}_{[11]}$ for the color configuration $\psi^{c}_{[2]}$ and $\psi^{osf}_{[2]}$ for $\psi^{c}_{[11]}$. All possible configurations are as follows:
$\psi^{c}_{[2]}\psi^{o}_{[2]}\psi^{s}_{[11]}\psi^{f}_{[2]},\ \psi^{c}_{[2]}\psi^{o}_{[2]}\psi^{s}_{[2]}\psi^{f}_{[11]},$
$\psi^{c}_{[2]}\psi^{o}_{[11]}\psi^{s}_{[11]}\psi^{f}_{[11]},\ \psi^{c}_{[2]}\psi^{o}_{[11]}\psi^{s}_{[2]}\psi^{f}_{[2]}$ (4)
for color sextet $\psi^{c}_{[2]}$, and
$\psi^{c}_{[11]}\psi^{o}_{[2]}\psi^{s}_{[11]}\psi^{f}_{[11]},\ \psi^{c}_{[11]}\psi^{o}_{[2]}\psi^{s}_{[2]}\psi^{f}_{[2]},$
$\psi^{c}_{[11]}\psi^{o}_{[11]}\psi^{s}_{[11]}\psi^{f}_{[2]},\ \psi^{c}_{[11]}\psi^{o}_{[11]}\psi^{s}_{[2]}\psi^{f}_{[11]}$ (5)
for color triplet $\psi^{c}_{[11]}$, where, $\psi^{o}_{[2]}$ and $\psi^{o}_{[11]}$ correspond to $l_{\chi_{1}}=0$ and $l_{\chi_{1}}=1$, respectively, as defined by the basis in Table 1. $\psi^{s}_{[2]}$ and $\psi^{s}_{[11]}$ correspond to spin $s_{qq}=1$ and $s_{qq}=0$, respectively. $\psi^{f}_{[2]}$ and $\psi^{f}_{[11]}$ correspond to isospin $I_{qq}=1$ and $I_{qq}=0$, respectively; analogous definitions apply to the antidiquark cluster. Thus, the configurations of total isospin $I=0$ and $I=1$ light tetraquarks are listed in Table 2.
Table 2: $I=0$ and $I=1$ Light tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$. Pairwise Pauli is enforced for $(12)$ and $(34)$.
Bases $(qq)$ $(\bar{q}\bar{q})$ Total II
$l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$} $\psi^{c}_{qq}$ $s_{qq}$ $I_{qq}$ $\psi^{c}_{\bar{q}\bar{q}}$ $s_{\bar{q}\bar{q}}$ $I_{\bar{q}\bar{q}}$
{1, 0, 0} $\bar{\mathbf{3}}_{c}$ 1 0 $\mathbf{3}_{c}$ 0 0 0
{0, 1, 0} $\mathbf{6}_{c}$ 1 0 $\bar{\mathbf{6}}_{c}$ 0 0 0
{0, 0, 1} $\mathbf{6}_{c}$ 1 0 $\bar{\mathbf{6}}_{c}$ 1 0 0
{0, 0, 1} $\bar{\mathbf{3}}_{c}$ 1 1 $\mathbf{3}_{c}$ 0 0 11
{0, 1, 0} $\bar{\mathbf{3}}_{c}$ 1 1 $\mathbf{3}_{c}$ 1 0 11
{1, 0, 0} $\bar{\mathbf{3}}_{c}$ 1 0 $\mathbf{3}_{c}$ 1 1 11
{0, 0, 1} $\mathbf{6}_{c}$ 1 0 $\bar{\mathbf{6}}_{c}$ 0 1 11
{0, 1, 0} $\mathbf{6}_{c}$ 1 0 $\bar{\mathbf{6}}_{c}$ 1 1 11
{1, 0, 0} $\mathbf{6}_{c}$ 1 1 $\bar{\mathbf{6}}_{c}$ 1 0 11
For $qq$ pair, color $\bar{\mathbf{3}}_{c}$ is antisymmetric (A), whereas $\mathbf{6}_{c}$ is symmetric (S). Analogous definitions for $\bar{q}\bar{q}$ pair, $\mathbf{3}_{c}$ and $\bar{\mathbf{6}}_{c}$ are A and S, respectively. Spin $s=0$ is A and $s=1$ is S. The internal spatial wave function is S for even $l{\chi}$ and A for odd $l{\chi}$. In $SU_{f}(2)$ flavor symmetry, $I=0$ and $I=1$ correspond to flavor A and S configurations, respectively. For the configuration in which both $I_{qq}=1$ and $I_{\bar{q}\bar{q}}=1$, the total isospin II can take the values 0, 11, or $2$, which may lead to ambiguous isospin eigenstates. This issue is not considered in this work.
For $cc\bar{c}\bar{c}$ tetraquarks, the $cc$ cluster has a symmetric flavor component $\psi^{f}_{[2]}$, so the possible configurations are as follows:
$\psi^{c}_{[2]}\psi^{o}_{[2]}\psi^{s}_{[11]}\psi^{f}_{[2]},\ \psi^{c}_{[2]}\psi^{o}_{[11]}\psi^{s}_{[2]}\psi^{f}_{[2]}$ (6)
for color sextet $\psi^{c}_{[2]}$,
$\psi^{c}_{[11]}\psi^{o}_{[2]}\psi^{s}_{[2]}\psi^{f}_{[2]},\ \psi^{c}_{[11]}\psi^{o}_{[11]}\psi^{s}_{[11]}\psi^{f}_{[2]}$ (7)
for color triplet $\psi^{c}_{[11]}$, and similarly for charm anti-diquark cluster. The possible configurations of fully-charm tetraquarks are listed in Table 3.
Table 3: Possible fully-charm tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
Bases $(cc)$ $(\bar{c}\bar{c})$
$l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$} $\psi^{c}_{cc}$ $s_{cc}$ $\psi^{c}_{\bar{c}\bar{c}}$ $s_{\bar{c}\bar{c}}$
{1, 0, 0} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 0
{0, 1, 0} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 0
{0, 0, 1} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 1
For $qc\bar{q}\bar{c}$ tetraquarks, there is no Pauli-antisymmetry constraint tying flavor to color, spin, or space, so no pair contains two identical quarks or two identical antiquarks. The color-spin-orbital (CSO) sector can therefore mix freely, independent of flavor, and charmonium-like tetraquark states are isospin-blind. In this work, we use the notation $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{\bar{3}\otimes 3}\psi^{s=1}_{(1\otimes 0)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 1)}$, $\psi^{c}_{6\otimes\bar{6}}\psi^{s=1}_{(1\otimes 0)}$ to denote these configurations. The possible configurations of charmonium-like tetraquark are listed in Table 4
Table 4: Charmonium-like tetraquark configurations for $J^{PC}=1^{-+}$ in a complete HO basis. Internal orbitals $l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}\in\{0,1\}$; $l_{\chi_{1}}+l_{\chi_{2}}+l_{\chi_{3}}=1$.
Bases $(qc)$ $(\bar{q}\bar{c})$
$l_{\chi_{1}},l_{\chi_{2}},l_{\chi_{3}}$} $\psi^{c}_{qq}$ $s_{qq}$ $\psi^{c}_{\bar{q}\bar{q}}$ $s_{\bar{q}\bar{q}}$
{0, 0, 1} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 0
{0, 1, 0} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 0
{1, 0, 0} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 0
{0, 0, 1} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 1
{0, 1, 0} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 1
{1, 0, 0} $\bar{\mathbf{3}}_{c}$ 1 $\mathbf{3}_{c}$ 1
{0, 0, 1} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 0
{0, 1, 0} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 0
{1, 0, 0} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 0
{0, 0, 1} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 1
{0, 1, 0} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 1
{1, 0, 0} $\mathbf{6}_{c}$ 1 $\bar{\mathbf{6}}_{c}$ 1
II.2 Hamiltonian
We employ a widely used nonrelativistic Hamiltonian Wang et al. (2021); Wu and Zhu (2025); Zhao et al. (2025) to study meson and tetraquark systems, which takes the following form:
$H=$ $H_{0}+H_{so},$
$H_{0}=$ $\sum_{k=1}^{N}(\frac{1}{2}M^{ave}_{k}+\frac{p_{k}^{2}}{2m{k}})$
$+\sum_{i<j}^{N}(-\frac{3}{16}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})(V_{0}(r_{ij})+V_{ss}(r_{ij})),$
$H_{so}=$ $\sum_{i<j}^{N}(-\frac{3}{16}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j})(V_{so}(r_{ij})),$ (8)
where $M^{ave}_{k}$ denotes the spin-averaged mass as $\frac{1}{4}M_{PS}+\frac{3}{4}M_{V}$. $m_{k}$ are the constituent quark masses. $\vec{\lambda}_{i}$ are the quark color operator in SU(3). The central potential $V_{0}(r_{ij})$, the spin-spin interaction $V_{ss}(r_{ij})$ and spin-orbital interaction $V_{so}(r_{ij})$ are taking the form:
$V_{0}$ $=A_{ij}r_{ij}-\frac{B_{ij}}{r_{ij}},$
$V_{ss}$ $=\frac{1}{6m_{i}m_{j}}\Delta V_{V}(r)\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$
$V_{so}$ $=\frac{1}{r_{ij}}\frac{dV_{V}}{d(r_{ij})}\frac{1}{4}\Bigg[\left(\frac{1}{m_{i}^{2}}+\frac{1}{m_{j}^{2}}+\frac{4}{m_{i}m_{j}}\right)\vec{L_{ij}}\cdot\vec{S_{ij}}$
$\quad+\left(\frac{1}{m_{i}^{2}}-\frac{1}{m_{j}^{2}}\right)\vec{L_{ij}}\cdot\left(\vec{s_{i}}-\vec{s_{j}}\right)\Bigg]$
$\quad-\frac{1}{r_{ij}}\frac{dV_{S}}{d(r_{ij})}\left(\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{2m_{i}^{2}}+\frac{\vec{L_{ij}}\cdot\vec{s_{i}}}{m_{j}^{2}}\right),$ (9)
where $\vec{\sigma}_{i}$ are the quark spin operator in SU(2). Note that we have employed $V_{V}(r)=-BErf[\sigma r]/r$ and $V_{S}(r)=Ar$, taken from Ref. Schoberl (1986). $m_{i}$ and $m_{j}$ are constituent quark masses of ith and jth quark. $\vec{s_{i}}$ represents the spin operator for ith quark. $\vec{S_{ij}}=\vec{s_{i}}+\vec{s_{j}}$ is the spin operator for (ij)th quark pair. $\vec{L_{ij}}$ is relative orbital operator, taking the form, $\vec{L_{ij}}=\vec{r_{ij}}\times\vec{p_{ij}}$, where $\vec{p_{ij}}=\frac{m_{i}\vec{p_{i}}-m_{j}\vec{p_{j}}}{m_{i}+m_{j}}$.
In line with the previous works Zhao et al. (2021, 2022, 2025), $A_{ij}$, $B_{ij}$, and $\sigma_{ij}$ are proposed to be mass dependent coupling parameters, taking the form
$A_{ij}=a+bm_{ij},\;\;B_{ij}=B_{0}\sqrt{\frac{1}{m_{ij}}},\;\;\sigma_{ij}=\sigma_{0}{m_{ij}}.$ (10)
with $m_{ij}$ being the reduced mass of $i$th and $j$th quarks, defined as $m_{ij}=\frac{2m_{i}m_{j}}{m_{i}+m_{j}}$. $a$, $b$, $B_{0}$, and $\sigma_{0}$ are constants.
The central potential $V_{0}$ and the spin-spin interaction $V_{ss}$ are treated as leading effects, while the spin-orbit interaction $V_{so}$ is regarded as a perturbation that shifts the spectrum. The Schrödinger equation is solved for mesons and tetraquarks with the Hamiltonian $H_{0}$, yielding the eigenvalue $E_{0}$ and eigenstates $\psi_{0}$. The mass spectrum is then obtained by adding the diagonalized $H_{so}$ in the basis of these eigenstates.
The coupling constants $a$, $b$, $B_{0}$, and $\sigma_{0}$, as well as the constituent quark masses $m_{u}$ and $m_{c}$, are taken from previous work Zhao et al. (2021, 2025). These parameters were fixed by reproducing the S- and P-wave mass spectra of light, charmed, bottom, charmonium, and bottomonium mesons within the Hamiltonian in Eq. (II.2). The imported parameters and constituent quark masses are listed in Table 5, and the corresponding fits can be found in Ref. Zhao et al. (2025).
Table 5: Imported model coupling parameters.
$a$ $b$ $B_{0}$ $\sigma_{0}$ $m_{u}$ $m_{c}$
[…]
II.3 Two-body strong decay
The strong two-body decay properties of the bound states, specifically the Y tetraquark states within the rearrangement mechanism, are studied, as shown in Figure 1.
$T_{csf}$ $=$ $\langle\psi_{f}^{csf}|\psi_{i}^{csf}\rangle$ (11)
$=$ $\langle\psi_{M_{1}M_{2}}^{c}|\psi_{Y}^{c}\rangle\langle\psi_{M_{1}M_{2}}^{s}|\psi_{Y}^{s}\rangle\langle\psi_{M_{1}M_{2}}^{f}|\psi_{Y}^{f}\rangle$
where $\psi_{Y}^{c}$, $\psi_{Y}^{s}$, and $\psi_{Y}^{f}$ are color, spin, and flavor wave functions of the initial states corresponding to the Y tetraquark states in mass spectrum. $\psi^{c}_{M_{1}M_{2}}$, $\psi^{s}_{M_{1}M_{2}}$ and $\psi^{f}_{M_{1}M_{2}}$ are the color, spin, and flavor wave functions of the final states of all possible two-body strong decay channels: $\omega h_{1}$ and $\eta f_{1}$ for isospin $I=0$ light tetraquarks; $\rho h_{1}$ and $\pi f_{1}$ for isospin $I=1$ light tetraquarks; $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ for charmonium-like tetraquarks; and $\eta_{c}\chi_{c1}$ and $J/\psi h_{c}$ for fully-charm tetraquarks.
In the work, the ratio between $|T_{csf}|^{2}$ is considered to be branching ratio approximately supposing that the spatial part from different decay channels provide similar contribution.
Figure 1: Quark rearranged diagram for $1^{-+}$ tetraquark.
III Results and Discussion
Table 6: Expectation values of hyperfine interaction of S=1 $qc\bar{q}\bar{c}$ $\langle\psi^{k1}_{c}\psi^{h1}_{s}|\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}|\psi^{k2}_{c}\psi^{h2}_{s}\rangle$
$\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)}\rangle$ $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 1)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)}\rangle$ $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 1)}\rangle$
$|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)}\rangle$ $\left(-\frac{8}{3},0,0,0,0,8\right)$ $\left(0,-\frac{4\sqrt{2}}{3},-\frac{4\sqrt{2}}{3},\frac{4\sqrt{2}}{3},\frac{4\sqrt{2}}{3},0\right)$ $\left(0,0,0,0,0,0\right)$ $\left(0,-4,4,-4,4,0\right)$
$|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 1)}\rangle$ $\left(0,-\frac{4\sqrt{2}}{3},-\frac{4\sqrt{2}}{3},\frac{4\sqrt{2}}{3},\frac{4\sqrt{2}}{3},0\right)$ $\left(-\frac{8}{3},\frac{4}{3},\frac{4}{3},\frac{4}{3},\frac{4}{3},-\frac{8}{3}\right)$ $\left(0,-4,4,-4,4,0\right)$ $\left(0,2\sqrt{2},-2\sqrt{2},-2\sqrt{2},2\sqrt{2},0\right)$
$|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)}\rangle$ $\left(0,0,0,0,0,0\right)$ $\left(0,-4,4,-4,4,0\right)$ $\left(-\frac{4}{3},0,0,0,0,-4\right)$ $\left(0,-\frac{10\sqrt{2}}{3},-\frac{10\sqrt{2}}{3},\frac{10\sqrt{2}}{3},\frac{10\sqrt{2}}{3},0\right)$
$|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 1)}\rangle$ $\left(0,-4,4,-4,4,0\right)$ $\left(0,2\sqrt{2},-2\sqrt{2},-2\sqrt{2},2\sqrt{2},0\right)$ $\left(0,-\frac{10\sqrt{2}}{3},-\frac{10\sqrt{2}}{3},\frac{10\sqrt{2}}{3},\frac{10\sqrt{2}}{3},0\right)$ $\left(\frac{4}{3},\frac{10}{3},\frac{10}{3},\frac{10}{3},\frac{10}{3},\frac{4}{3}\right)$
The color operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}$ in the Hamiltonian in Eq. (II.2), along with the mass dependent coefficient in the Cornell-like potential, mixes different color configurations, leading to nonzero cross terms for tetraquarks.
Similarly, the color-spin operator $\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$, together with the mass dependent coefficient in the spin-spin interaction,
$\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}V_{ss}$ $=\frac{2B_{ij}\sigma_{ij}^{3}}{3m_{i}m_{j}\sqrt{\pi}}e^{-\sigma_{ij}^{2}r_{ij}^{2}}\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j},$ (12)
mixes different color-spin configurations.
For reference, the expectation values of $O_{ij}=\vec{\lambda}_{i}\cdot\vec{\lambda}_{j}\vec{\sigma}_{i}\cdot\vec{\sigma}_{j}$ for $S=1$ states are listed in Table 6, with all components given in the order $(O_{12},O_{13},O_{23},O_{14},O_{24},O_{34})$. The mass dependent factor in Eq. (12) induces nonzero cross terms between different color-spin configurations.
In conclusion, the $I=0$ light tetraquarks are linear combinations of $|\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 1)}\rangle$, and $I=1$ light tetraquarks are linear combinations of $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 1)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 1)}\rangle$. For fully-charm tetraquark, $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 1)}\rangle$ configurations are considered due to the quark configurations we have mentioned in Sec. II.1. The charmonium-like tetraquarks are linear combinations of $|\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 0)},$ $\psi^{c}_{6\otimes\bar{6}}\psi^{S=1}_{(1\otimes 1)},$ $\psi^{c}_{\bar{3}\otimes 3}\psi^{S=1}_{(1\otimes 1)}\rangle$.
After testing the convergence of the mass calculation for the three tetraquark configurations, light ($qq\bar{q}\bar{q}$), charmonium-like ($qc\bar{q}\bar{c}$), and fully-charmed ($cc\bar{c}\bar{c}$), we use $n=6$ as the principal quantum number for all cases. The choice yields stable results and corresponds to a basis size of $N=13$. The masses obtained with this setting are taken as the final results in this section.
III.1 Results of light tetraquark
The mass results for the mixed light tetraquark states are given in table 7, where En (n = 1,2,3,…) represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) with color-spin configuration mixing. The branching ratios of the E1 to E6 with various mixed configurations are also listed in Tables 7. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
Table 7: Masses for mixed $qq\bar{q}\bar{q}$ tetraquark. The mass results are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
$qq\bar{q}\bar{q}$ THR E1 E2 E3 E4 E5
$M_{cal}$ $I=0$ 1953 2210 2233 2307 2504
$\eta f_{1}(1285)$ 1835 2 <1<1 1 54 16
$\omega h_{1}(1170)$ 1952 <1<1 12 273 117
$M_{cal}$ $I=1$ 1880 1920 2104 2239 2264
$\pi f_{1}(1285)$ 1420 49 44 1 <1<1 <1<1
$\rho h_{1}(1170)$ 1940 1 1 1
For an S-wave decay of an exotic $J^{PC}=1^{-+}$ state, the only two-meson combinations allowed by the quantum numbers are $\left(0^{-+}\oplus 1^{++}\right)$ and $\left(1^{--}\oplus 1^{+-}\right)$. All other $S+P$ (S-wave + P-wave meson) combinations either fail to satisfy $C=+$ or cannot produce $J=1$ when the relative orbital angular momentum between the two mesons is $L_{rel}=0$.
It is worth mentioning that, in principle, the $I=0$ $1^{-+}$ tetraquark can decay through the $\eta\eta^{\prime}$ channel when the relative orbital angular momentum is $L_{rel}=1$, which allows $J=1$ and negative parity $P=$-. For a two-meson final state composed of two different self-conjugate mesons $M_{1}\neq M_{2}$ (each with a well-defined intrinsic $C_{M_{1}},\,C_{M_{2}}=\pm 1$), the total two-body C-parity is simply $C_{tot}=C_{M_{1}}C_{M_{2}}$, because there is no interchange symmetry between the two distinct mesons. Thus, for the $\eta\eta^{\prime}$ channel, the charge conjugation can take $C_{tot}=C_{M_{1}}C_{M_{2}}=+$ Shi et al. (2024a); Matveev et al. (2025).
However, in the work, the spin-color-flavor factor of the transition amplitude is zero in the two-body strong decay to $\eta\eta^{\prime}$ from $I=0$ $1^{-+}$ light tetraquarks.
III.2 $\pi_{1}(2015)$
E852 observed two $1^{-+}$ states while studying $\pi^{-}p$ collisions in the decay channel $f_{1}(1285)\pi$ Kuhn et al. (2004) around 1.6 and 2.0 GeV respectively. The higher-mass peak is as $M=2001\pm 30\pm 92~\mathrm{MeV}$ and $\Gamma=333\pm 52\pm 49~\mathrm{MeV}$, commonly denoted $\pi_{1}(2015)$ (a.k.a. $\pi_{1}(2000)$). Later, the $\pi_{1}(2015)$ state is observed again by E852 in the decay channel $b_{1}(1235)\pi$ Lu et al. (2005), with mass and decay width $M=2014\pm 20\pm 16~\mathrm{MeV}$ and $\Gamma=230\pm 32\pm 73~\mathrm{MeV}$ which are consistent with that observed in the $f_{1}(1285)\pi$ final state.
VES also performs the similar partial-wave analysis, and the $\pi_{1}(1600)$ is observed in both $f_{1}(1285)\pi$ and $b_{1}(1235)\pi$ decay channels Amelin et al. (2005). Their intensity of the $1^{-+}$ partial wave in both channel above 1.9 GeV is considered to be consistent with that observed by E852 Meyer and Swanson (2015). However, VES did not comment on this or claimed the existence of the $\pi_{1}(2015)$.
COMPASS recently presents a prelimnary results of partial-wave decomposition of the diffractive reaction $\pi^{-}p\to\pi^{-}\pi^{+}\pi^{-}\eta p$ Spülbeck (2025), and no clear evidence for an excited state as previously reported by E852 Kuhn et al. (2004) is found, and they highlight the model dependence of such interpretations.
A recent light hybrid nonet calculation in a semirelativistic quark potential model organizes the light $1^{-+}$ $\pi_{1}(1600)$ states into a hybrid nonet with 1669 MeV correspond to the hybrid configurations $(u\bar{u}-d\bar{d})g/\sqrt{2}$. The following analyses accommodate $\pi_{1}(2015)$ to be first radial excitation of $\pi_{1}(1600)$ near $2.1~\text{GeV}$ Chen and Liu (2025). In this view, $\pi_{1}(2015)$ may have sizable hybrid content but requires confirmation in $f_{1}\pi$ and $b_{1}\pi$ with modern amplitude methods.
[…]
The observed S+PS{+}P decay pattern matches expectations for hybrids, yet is also natural for a compact tetraquark. a QCD sum rule analysis of exotic tetraquark states, using diquark-antidiquark currents with $I^{G}J^{PC}=1^{-}1^{-+}$ and $0^{+}1^{-+}$Chen et al. (2008a, b) can accommodate $\pi_{1}(2015)$, with the $qs\bar{q}\bar{s}$ configuration around $2.0~\text{GeV}$, which supports the assignment of $\pi_{1}(2015)$ to a compact tetraquark that naturally prefers S+PS{+}P decays such as $b_{1}\pi$ and $f_{1}\pi$.
In this work, the $I=1$ light tetraquark E2 state with a mass of around 2 GeV is comparable to the $\pi_{1}(2015)$ within the experimental uncertainties. The $\pi_{1}(2015)$ may be assigned to the E2 tetraquark state based on its large decay ratio to $\pi f_{1}$. We suggest searching for $I=1$ tetraquark states in $S+P$ channels around 1.9–2.0 GeV, and for $I=0$ tetraquark states in $S+P$ channels around 2.3 GeV and 2.5 GeV, motivated by the large decay ratios predicted for several tetraquark states.
III.3 $\eta_{1}(1855)$
In 2022, the BESIII Collaboration reported the discovery of a new isoscalar exotic state, $\eta_{1}(1855)$ with quantum numbers $J^{PC}=1^{-+}$, observed in the decay process $J/\psi\to\gamma\eta\eta^{\prime}$ Ablikim et al. (2022b). The measured mass is $(1855\pm 9^{+6}_{-1})$ MeV, and the width is $(188\pm 18^{+3}_{-8})$ MeV, with a statistical significance greater than $19\sigma$.
BESIII also estimated the product branching fraction as $\mathcal{B}(J/\psi\to\gamma\eta_{1}(1855))\times\mathcal{B}(\eta_{1}(1855)\to\eta\eta^{\prime})=(2.70\pm 0.41^{+0.16}_{-0.35})\times 10^{-6}$ Ablikim et al. (2022b). This state is considered a candidate for a hybrid meson, in agreement with lattice QCD predictions Dudek et al. (2013).
In the hybrid meson framework, the observation of $\eta_{1}(1855)$, together with the isovector $\pi_{1}(1600)$, provides valuable information for understanding the $1^{-+}$ hybrid nonet Qiu and Zhao (2022); Chen et al. (2022); Shastry et al. (2022); Wang et al. (2022); Chen et al. (2023a); Swanson (2023); Chen et al. (2023b); Shastry and Giacosa (2023). A semirelativistic quark potential model was used to estimate the masses of light hybrid mesons with $J^{PC}=1^{-+}$ Chen et al. (2023b). This study suggests that $\pi_{1}(1600)$ and $\eta_{1}(1855)$ might correspond to the hybrid configurations $(u\bar{u}-d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$, respectively.
However, another recent study modeled hybrid mesons as three-body systems (quark–antiquark–gluon), treating the gluon as a real constituent Ma et al. (2025). Their results indicate that both $\pi_{1}(1600)$ and $\eta_{1}(1855)$ cannot be interpreted as $1^{-+}$ hybrids at the same time, due to inconsistencies in their decay widths. In particular, they found that the decay width of $\eta_{1}(1855)\to\eta\eta^{\prime}$ is nearly zero at leading order, which challenges its hybrid interpretation.
Alternatively, in the hadronic molecule picture, $\eta_{1}(1855)$ could be a bound state of $K\bar{K}_{1}(1400)$, as its mass is about 40 MeV below this threshold Dong et al. (2022); Yang et al. (2023). Ref. Dong et al. (2022) used the one-boson exchange model to explore the binding and decay properties of such a molecule and found results supporting this interpretation. Similarly, Ref. Yang et al. (2023) studied its radiative and strong decays using the effective Lagrangian approach and compared them with experimental data.
On the other hand, a study using the chiral quark model did not support the $K\bar{K}_{1}$ molecular interpretation of $\eta_{1}(1855)$ Tan et al. (2025). In another analysis using QCD sum rules Liu et al. (2025), the exotic $K\bar{K}_{1}(1270)$ and $K\bar{K}_{1}(1400)$ molecular states with $J^{PC}=1^{-+}$ were investigated, but the spectral functions turned out negative in physical regions, making the results unreliable.
A tetraquark molecular picture was also considered. In Ref. Wan et al. (2022), configurations like $1_{c}]s\bar{s}\otimes[1_{c}]q\bar{q}$ and $1_{c}]s\bar{q}\otimes[1_{c}]s\bar{q}$ were studied using QCD sum rules. The mass calculations suggest that $\eta_{1}(1855)$ could match the $1_{c}]s\bar{s}\otimes[1_{c}]q\bar{q}$ configuration.
In the work, the theoretical results of branching ratio for $1^{-+}$ light tetraquark state decaying to $\eta\eta^{\prime}$ are all equal to zero. Thus, the $\eta_{1}(1855)$ state observed from $\eta\eta^{\prime}$ Ablikim et al. (2022b) cannot be accommodated in the current tetraquark picture, and the state could be other structures, for example, a hadronic molecule state, or a hybrid meson as reviewed above. While some models support its exotic structure, others challenge these assignments, highlighting the need for further theoretical and experimental studies.
III.4 Results of charmonium-like $1^{-+}$ tetarquark
Lattice QCD predicts that the lightest $1^{-+}$ charmonium hybrid ($\eta_{c1}$) lies around $4.2-4.4~\mathrm{GeV}$ and decays mainly to open-charm $S+P$ channels such as $_{1}\bar{D}, DD$$\bar{D}$, and $D\bar{D}^{*}$. Its hidden-charm branching fractions are small, which makes it hard to observe in clean but suppressed radiative modes Shi et al. (2024b).
At present, no charmonium-like state with $J^{PC}=1^{-+}$ has been confirmed. Extensive studies of the $XYZ$ sector by BESIII, LHCb, Belle, and Belle II have found many states with non-exotic, non-$1^{-+}$ quantum numbers (e.g. $1^{++},1^{+-},0^{++},2^{++}$), while searches for $1^{-+}$ have so far found nothing. A dedicated BESIII search, $e^{+}e^{-}\to\gamma\eta^{(\prime)}\eta_{c}$ at $\sqrt{s}=4.258\text{-- }4.681~\mathrm{GeV}$, saw no significant signal and set 90% C.L. upper limits on the cross sections Ablikim et al. (2025).
Promising search modes that act as quantum-number filters for $C=+$, odd parity, and $J=1$ are: (i) radiative production at BESIII/Belle II, $e^{+}e^{-}\to\gamma Y(1^{-+})$ with $Y\to S{+}P$ meson pairs ($\eta+\chi_{cJ}$, $\rho+h_{c}$), which provide clean angular observables for amplitude fits Chao et al. (2013); Zhang et al. (2025); and (ii) $B$-meson decays at LHCb/Belle II, $B\to KY(1^{-+})$ with $Y\to S{+}P$ meson pairs, enabling direct tests of the exotic $1^{-+}$ hypothesis.
Higher-luminosity scans and better Electromagnetic Calorimeter (EMC) calibrations at BESIII (below 4.66 GeV) and Belle II (10.58 GeV), plus targeted angular-moment analyses for $1^{-+}$ are especially promising Yuan (2021). Notably, BESIII’s first direct production of a $C=+$ state, $e^{+}e^{-}\to\chi_{c1}$, shows that $e^{+}e^{-}$ experiments can access $C=+$ states Ablikim et al. (2022c). The work confirms key tools, photon reconstruction and interference control required for a $1^{-+}$ search.
The mass spectrum and S-wave decay ratio are listed in Table 8 for the charmonium-like tetraquark. The lowest mass is around 4.2 GeV with a relative large ratio to $\eta\chi_{c1}$ channel. In the work, the theoretical results of branching ratio for $1^{-+}$ charmonium-like tetraquark state decaying to $\eta\eta_{c}$ are very small. The work suggests searching for tetraquark states in $\eta\chi_{c1}$ channels in the mass range of 4.24.34.2-4.3 GeV due to the large decay ratio of two tetraquark states.
Table 8: Masses for mixed tetraquark. The theoretical mass results $M_{cal}$ are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
$qc\bar{q}\bar{c}$ THR E1 E2 E3 E4 E5 E6
$M_{cal}$ 4237 4279 4281 4300 4415 4435
$\rho h_{c}$ 4295 5 <<1 1
$\eta\chi_{c1}$ 4060 18 24 1 4 <<1 <<1
III.5 Results of fully-charm $1^{-+}$ tetarquark
To date there is no confirmed observation of a $1^{-+}$ fully-charm state. The established structures in the di-$J/\psi$ spectrum, most prominently the narrow $X(6900)$ and companion peaks near $6.6$ and $7.1$ GeV observed by LHCb Aaij et al. (2020), CMS Hayrapetyan et al. (2024), and ATLAS Aad et al. (2023), are studied in the previous work Zhao et al. (2021), and tentatively assigned as S-wave tetraquark states. A recent CMS angular analysis determines the parity and charge-conjugation of the fully-charm peaks to be $2^{++}$ and disfavors $J=1$ assignments, thus excluding $1^{-+}$ for those structures Hayrapetyan et al. (2025). Some theoretical studies typically locate low-lying fully-charm tetraquarks with $0^{++}$, $2^{++}$ (and sometimes $1^{+-}$), not $1^{-+}$ Karliner and Rosner (2020); Yang et al. (2021).
The most sensitive venues are CMS/ATLAS and LHCb, which provide large prompt-charm samples and clean muon triggers for di-charmonium studies. However, the main channel $J/\psi J/\psi$ favors even-$J$ and $C=+$ waves; $J=1$ in S-wave is strongly suppressed near threshold and disfavored by the Bose symmetry of two identical vectors, so sensitivity to $1^{-+}$ is low. Inclusive production also dilutes signals with large non-resonant di-charmonium backgrounds, and limited data in $S+P$ modes hinder full amplitude analyses. A better strategy is to move beyond identical-vector final states to $S+P$ charmonium pairs with definite $C=+$ and richer angular structure (e.g. $\eta_{c}\chi_{cJ}$, $J/\psi h_{c}$).
On the detector side, LHC Run 3 upgrades, higher-rate di-muon triggers, better mass resolution, and dedicated di-charmonium selections, should increase sensitivity Hayrapetyan et al. (2024); Aad et al. (2025). At the $\bar{p}p$ threshold, PANDA will enable formation scans with excellent resolution and access to exclusive $S+P$ final states, which could be decisive for exotic $J^{PC}$ tests Pelizaeus (2012). Belle II can also contribute via double-charmonium production at 10.58 GeV, with full-amplitude analyses in $J/\psi h_{c}$ and $\eta_{c}\chi_{cJ}$, though cross sections are lower than at the LHC.
Table 9: Masses for $1^{-+}$ $cc\bar{c}\bar{c}$ tetraquark. The mass results are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
$cc\bar{c}\bar{c}$ THR E1 E2 E3 E4 E5 E6
$M_{cal}$ 6586 6754 6870 6903 7059 7074
$J/\psi h_{c}$ 6622 1 10 310 247 1
$\eta_{c}\chi_{c1}$ 6496 54 1 10 310 247 1
The mass spectra and S-wave decay ratio are listed in Tables 9 for the fully-charmed tetraquark. The lowest mass is around 6.6 GeV with a relative large ratio to $\eta_{c}\chi_{c1}$ channel. The work suggests searching for tetraquark states in $\eta_{c}\chi_{c1}$ channels around 6.6 GeV, and in $J/\psi h_{c}$ and $\eta_{c}\chi_{c1}$ channels in the mass range of 6.9-7.1 GeV, due to the large decay ratio of several tetraquark states.
IV Summary
While the hybrid interpretation has been extensively explored in the literature, the tetraquark scenario remains particularly compelling due to the rich internal dynamics inherent to four-quark systems.
The masses and decay branching ratio of $1^{-+}$ light, charmonium-like, and fully-charm tetraquark states are studied using a constituent quark model (CQM) that incorporates a Cornell-like potential and the Breit-Fermi interaction. The theoretical results for tetraquarks are compared with those of $1^{-+}$ exotic states.
The lowest $1^{-+}$ light tetraquark state is around 1.9 GeV, which is significantly far from the observed $\pi_{1}(1400)$ and $\pi_{1}(1600)$ states.
The observed states $\eta_{1}(1855)$ and $\pi_{1}(2015)$ are located within the light tetraquark range. The theoretical results of branching ratio for $1^{-+}$ Y tetraquark state decaying to $\eta\eta^{\prime}$ are equal to zero. Thus, the $\eta_{1}(1855)$ state observed from $\eta\eta^{\prime}$ by BESIII cannot be accommodated in the current tetraquark picture.
The $\pi_{1}(2015)$ is assigned as $I=1$ E2 $1^{-+}$ light tetraquark state according to both mass agreement and the large decay ratio. More experimental data and theoretical studies are essential for making unambiguous assignments.
For the $1^{-+}$ charmonium-like tetraquark state, the lowest mass is predicted near 4.2 GeV and could decay into $\eta\chi_{c1}$ channel.
The lowest $1^{-+}$ fully-charmed tetraquark state is around 6.6 GeV and may be searched in $\eta_{c}\chi_{c1}$ channel.
Acknowledgements.
This work was supported by (i) Suranaree University of Technology (SUT), (ii) Thailand Science Research and Innovation (TSRI), and (iii) National Science Research and Innovation Fund (NSRF), Project No. 179349.